OpenStudy (anonymous):

What is the solution set of |x - 6| = 8? What is the solution set of |x + 3| - 2 = -5?

OpenStudy (anonymous):

two different questions?

OpenStudy (anonymous):

yeah

OpenStudy (linknissan):

x=8+6 or X=-8+6

OpenStudy (anonymous):

explain your answer don't just post.

OpenStudy (anonymous):

\[|x-6|=8\]\[x-6=\pm8\]\[x=\pm8+6....x=-2..or..x=14\] \[|x+3|-2=-5\]\[|x+3|=-3\]\[x+3=\pm3\]\[x=\pm3-3....x=-6...or...x=0\]

OpenStudy (linknissan):

its understood very simple..the person who posted this question dont undst..inbox me..

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct